The present age of a woman is 30years older than her daughter.15years ago, she was twice as old as her daughter.How old is her daughter at present.How old would the woman be 12years from now?

Answer by stanbon(75887)

The present age of a woman is 30years older than her daughter.15years ago, she was twice as old as her daughter.How old is her daughter at present.
Equations:
w = d + 30
(w-15) = 2(d-15)
----
Substitute for "w" and solve for "d":
d+30-15 = 2d-30
d = 45 yrs (daughter's age now)
---
Solve for "w";
w = d + 30
w = 45+30 = 75 yrs (woman(mother)age now)
========================
How old would the woman be 12years from now?
Ans: 75+12 = 87 yrs
========================
Cheers,
Stan H.
==============